Thursday, July 15, 2010

Summary Night Report 20100715

Report for night 2010-07-15

SA: Encarni
SO: Patrick
Others: Deon, Morgan, Hitesh, the IQ team, Chris, Zolisa, Siphelo

Notes and messages

* Did not open up tonight because of the weather. SAMS tests all night.

* EDS external weather conditions got stuck. BMS needed restarting to
get them back.

No comments:

Post a Comment